Aug 19, 2010 Calorimetry
Teleconference Time: 11:00 a.m. EDT
- ESNET and EVO session
- Phone in CC F326 is 757-269-6337
- Phone in the Regina Videoconference Suite is 306-585-4204
- Phone only if requested:
- +1-800-377-8846 : US
- +1-888-276-7715 : Canada
- +1-302-709-8424 : International
- then enter participant code: 77438230# (remember the "#"). Phone line by request only.
Items for followup from previous meeting(s)
- Elton and Ashot to meet and resolve any discrepancies between the simulations.
- Report cosmic-ray normalization of energy conversion to photons.
- SiPM evaluations
- UofR: Complete cosmic-ray energy calibration.
- IU: a) take into account 5 times lower rate in upstream compared to downstream b) geometry for summed units
- JLab: Complete testing of radiation damage to sensors, temperature dependence, etc.
- Review of Action Items
- Construction status
- To shashlyk or not to shashlyk.
- Construction Status
- Cosmic-ray normalization runs
- Other Updates
Attendees: Beni, Eugene, John, Fernando, Yi, Tim (JLab), Will (USM), Andrei, Zisis (Regina)
- Announcements: none
- Review of Action Items
- Monitoring: Athens will conduct tests on the USM light guides before the Collaboration Meeting and report then.
- FCAL simulations: deferred until Elton, Matt and Ashot attend (next?) video conference.
- SiPM Evaluations: Yi reported that Hamamatsu was contacted and will ship 12 1x1mm^2 asap (in 2-3 weeks). Measurements on single Photonique 2x2mm^2 tagger SiPM are under way. Noise is < 10 MHz; 1x1mm^2 typically are at 1MHz. Peltier and 0 degree measurements are under way. Hall-A radiation is 2.3 times that of AmBe but noise ratio after self-annealing is 4.5. Eugene reported that Carl's tests on the outlier SiPM show a factor 2-3 worse in noise, but all other parameters are the same as the other SiPMs.
- Calibration: see above, under Action Items
- Construction status: no report
- Eugene reported that Moscow Company will ship PMTs on August 25th to IU.
- Fernando reported that P. Smith has finished the redesign of the bases and Fernando is reviewing it. We will invite Paul to join us in two weeks together.
- To shashlyk or not to shashlyk: deferred, as mentioned above.
- Construction Status: 3rd shipment arrived at JLab, so 12 modules are now at Warwick storage facility. Mods 13-16 are being machined, and should finish by end of August so that the 4th shipment will leave for JLab in early Sept. Epoxy consumption for Mods 15-16 was 10-15% lower than average, and we will increase it somewhat (as an extra a safety margin). Standing order from Bicron for epoxy has been implemented, with 10 gallons arriving at the middle of each month, with 30 gallon reserve. All else is nominal.
- Delamination tests were carried out for lead samples bonded to machined aluminum pieces using industrial epoxy, for a) normal cleaning with ethanol and wipes, and b) 5% white vinegar, then ethanol and wipes; results: 11000 N vs 22000 N, so method b) looks very promising. Discussion ensued and a few more checks will be done, to ensure that the vinegar does not affect the optical epoxy, etc.
- Zisis will prepare statistics on the construction speed and present them at the collaboration meeting.
- Cosmic-ray normalization runs: we are collecting more statistics at the middle of the module for cross checks. We hope to be able to report at the next meeting, but more likely at the collaboration meeting the week after.
- Simulations: no report.
- Other Updates: Zisis reminded people to register for the collaboration meeting, whether they are coming or not. In two weeks, during the next Calorimetry meeting (Sept 2) we will finalize the talks to be presented at the meeting.